# The Gauge Theory Bootstrap: Predicting pion dynamics from QCD

### v2

@@ Appendix B Parameterization of the spectral density
In this appendix we describe the parameterizations used for the spectral density, designed to satisfy the correct low energy behavior as described in section 2.1. Consider, for example the vector current with the behavior (2.2b) at lowest order in $\chi_{PT}$ [41] in our conventions. Setting $m_{\pi}=1$, for $s\in\mathbb{R}_{\geq 4}$:
$\rho^{1}_{1}(s)=\frac{1}{(2\pi)^{4}}\frac{s}{24\pi}\left(1-\frac{4}{s}\right)^{\frac{3}{2}}=\frac{1}{(2\pi)^{4}}\frac{1}{192\pi}(s-4)^{\frac{3}{2}}\left[1-\frac{3}{8}(s-4)+\frac{15}{128}(s-4)^{2}+\ldots\right]$ (B.16)
@@
$\rho^{1}_{1}(s)=\frac{1}{(2\pi)^{4}}\frac{1}{192\pi}(s-4)^{\frac{3}{2}}\left[B+\sum_{n\geq 0}b_{n}(s-4)^{n}\right]$ (B.17)
@@
$\rho_{0}^{0}$ $=$ $-\frac{1}{(2\pi)^{4}}\frac{3}{64\pi}\left(1-\frac{s_{a}}{4}\right)^{\frac{1}{2}}\mathrm{Im}\left\{(1-z)\left[A+(1-z)\sum_{n=0}^{N}a_{n}z^{n}\right]\right\}$ (B.20)
$\rho_{1}^{1}$ $=$ $+\frac{1}{(2\pi)^{4}}\frac{1}{96\pi}\left(1-\frac{s_{a}}{4}\right)^{\frac{3}{2}}\mathrm{Im}\left\{(1-z)^{3}\left[B+(1-z)\sum_{n=0}^{N}b_{n}z^{n}\right]\right\}$ (B.21)
$\rho_{2}^{0}$ $=$ $-\frac{1}{(2\pi)^{4}}\frac{1}{640\pi}\left(1-\frac{s_{a}}{4}\right)^{\frac{5}{2}}\mathrm{Im}\left\{(1-z)^{5}\left[C+(1-z)\sum_{n=0}^{N}c_{n}z^{n}\right]\right\}$ (B.22)
@@ Appendix C Finite energy sum rules up to 3-loop
Finally, for the $S0$ and $P1$ waves, we find useful to improve the procedure by using a different analytical kernel such as the one in [68] that captures more on the lower energy region 111111as compared with the weight $s^{n}$ that tends to favor the larger values of $x$ close to $s_{0}$:
@@
In the main text we use these “pinched” sum rules for $\rho^{0}_{0},\rho^{1}_{1}$ up to $n=6$. The corresponding numerical coefficients $c_{i,j}$ are shown in tables 5, 6 for reference.
@@
Table 5: Numerical values of the S0 sum rule coefficients $c^{S0}_{i,j}$.

### v3

@@ Appendix B Parameterization of the spectral density
In this appendix we describe the parameterizations used for the spectral density, designed to satisfy the correct low energy behavior as described in section 2.1. Consider, for example the vector current with the behavior (2.2b) at lowest order in $\chi_{PT}$ [42] in our conventions. Setting $m_{\pi}=1$ we obtain, for $s\in\mathbb{R}_{\geq 4}$:
$\rho^{1}_{1}(s)=\frac{1}{(2\pi)^{4}}\frac{s}{24\pi}\left(1-\frac{4}{s}\right)^{\frac{3}{2}}=\frac{1}{(2\pi)^{4}}\frac{1}{48\pi}(s-4)^{\frac{3}{2}}\left[1-\frac{1}{8}(s-4)+\frac{3}{128}(s-4)^{2}+\ldots\right]$ (B.16)
@@
$\rho^{1}_{1}(s)=\frac{1}{(2\pi)^{4}}\frac{1}{48\pi}(s-4)^{\frac{3}{2}}\left[B+\sum_{n\geq 0}b_{n}(s-4)^{n}\right]$ (B.17)
@@
$\rho_{0}^{0}$ $=$ $-\frac{1}{(2\pi)^{4}}\frac{3}{32\pi}\left(1-\frac{s_{a}}{4}\right)^{\frac{1}{2}}\mathrm{Im}\left\{(1-z)\left[A+(1-z)\sum_{n=0}^{N}a_{n}z^{n}\right]\right\}$ (B.20)
$\rho_{1}^{1}$ $=$ $+\frac{1}{(2\pi)^{4}}\frac{1}{48\pi}\left(1-\frac{s_{a}}{4}\right)^{\frac{3}{2}}\mathrm{Im}\left\{(1-z)^{3}\left[B+(1-z)\sum_{n=0}^{N}b_{n}z^{n}\right]\right\}$ (B.21)
$\rho_{2}^{0}$ $=$ $-\frac{1}{(2\pi)^{4}}\frac{1}{320\pi}\left(1-\frac{s_{a}}{4}\right)^{\frac{5}{2}}\mathrm{Im}\left\{(1-z)^{5}\left[C+(1-z)\sum_{n=0}^{N}c_{n}z^{n}\right]\right\}$ (B.22)
@@ Appendix C Finite energy sum rules up to 3-loop
Finally, for the $S0$ and $P1$ waves, we find useful to improve the procedure by using a different analytical kernel such as the one in [69] that captures more on the lower energy region 121212as compared with the weight $s^{n}$ that tends to favor the larger values of $x$ close to $s_{0}$:
@@
In the main text we use these “pinched” sum rules for $\rho^{0}_{0},\rho^{1}_{1}$ up to $n=6$. (For $\rho^{0}_{2}$ we use the same sum rules with weights $s^{n}$ as in previous work [3].) The corresponding numerical coefficients $c_{i,j}$ are shown in tables 5, 6 for reference. The numerical values are used to constrain the spectral densities following eq.(C.28).
@@
Table 5: Numerical values of the S0 sum rule coefficients $c^{S0}_{i,j}$. These coefficients are used to constrain the spectral density